# Preference Elicitation for Offline Reinforcement Learning

### v1

@@ 7 Experimental results
We also measure how the coverage of the optimal policy affects performance in our setting. In Figure 3(b), we vary the behavioral policy $\pi_{\beta}$ underlying the offline data, ranging from optimal (density ratio coefficient $=1$) to highly suboptimal (large density ratio coefficient). The concentrability terms $C_{T}$ and $C_{R}$ are challenging to measure as they require considering entire function classes, but we report the accuracy of the maximum likelihood estimate for both models in Appendix D. We observe that preference elicitation methods perform best when the data is close to optimal (with the exception of a fully optimal, non-diverse dataset making reward learning from preferences challenging). More preference samples are required if the observational dataset has poor coverage of the optimal policy (large $C_{T}(\mathcal{F}_{T},\pi^{*})$), as the transition and reward models become less accurate for the trajectory distribution of interest.
@@
(b) As a function of optimal coverage.
Figure 3: Preference sample complexity $N_{p}$ as function of the properties of the observational data, to reach a suboptimality gap of $\epsilon=20$ over normalized environment returns (Star MDP). Mean and 95% confidence intervals over 20 experiments. $\times$ marks when the target suboptimality could not be achieved. In (b), we vary the density ratio coefficient – an upper bound for $C_{T}(\mathcal{F}_{T},\pi^{*})$ since it is challenging to quantify. See Appendix D for accuracy of the transition and reward models.
Gridworld and Sepsis Simulation. Finally, we validate our empirical findings on more complex environments detailed in Appendix C. We explore a gridworld experiment as well as a simulation of patient evolution in the intensive care unit from Oberst and Sontag [2019], designed to model the evolution of patients with sepsis. This medically-motivated example highlights another interesting and potentially important advantage of Sim-OPRL over OPRL, as it does not require feedback on real trajectories from the observational dataset. In a sensitive setting such as healthcare where data access is carefully controlled, it may be attractive to query experts about synthetic trajectories rather than real data samples. Sample complexity results across the different environments considered are given in Table 2, with similar conclusions: Sim-OPRL affords a higher preference sampling efficiency than OPRL baselines. For the sepsis environment, we note the number of preference samples needed to achieve our target suboptimality is very large, likely due to the sparse nature of the reward function. In a real-world application, we could potentially warm-start the reward model by leveraging proxy rewards signals in the offline data [Yu et al., 2021].
Table 2: Comparison of preference sample complexity $N_{p}$ with different sampling methods, to reach a suboptimality gap of $\epsilon=20$ over normalized returns. Mean and 95% confidence interval over 20 experiments. The best-performing offline method is highlighted in bold. Note that PbOP has an advantage by having access to direct interaction with the environment.
Star MDP (Figure 1) 32 $\pm$ 4 30 $\pm$ 4 4 $\pm$ 2 4 $\pm$ 2
Gridworld 105 $\pm$ 11 66 $\pm$ 7 49 $\pm$ 7 32 $\pm$ 4
Sepsis Simulation 18,856 $\pm$ 427 2,246 $\pm$ 143 830 $\pm$ 88 261 $\pm$ 59
@@ 8 Conclusion
We provide theoretical guarantees on the sample complexity of both approaches, demonstrating that performance depends on how well the offline data covers the optimal policy. Empirical evaluations in various environments confirm the practical effectiveness of our algorithm, as Sim-OPRL consistently outperforms OPRL baselines across different environments.
@@ A.1 Maximum Likelihood and Confidence Intervals
$\beta=\frac{c_{MLE}\log(\mathcal{N}_{\mathcal{F}_{g}}(1/N)/\delta)}{N},$
@@
Let $\bar{\mathcal{B}}$ be a $1/N$-bracket of $\mathcal{F}_{g}$ with |¯$|\bar{\mathcal{B}}|=\mathcal{N}_{\mathcal{F}_{g}}(1/N)$. Denote the set of all right brackets in $\bar{\mathcal{B}}$ by ~=$\tilde{\mathcal{B}}=\{b:\exists b^{\prime}s.t.[b^{\prime},b]\in\bar{\mathcal{B}}\}$. For $b\in\tilde{\mathcal{B}}$, we have:
@@
$P\left(\sum_{n=1}^{N}\log\frac{b(y^{n}|x^{n})}{g(y^{n}|x^{n})}>\log(1\delta)\right)$ $\leq\mathbb{E}\left[\exp\left(\sum_{n=1}^{N}\log\frac{b(y^{n}|x^{n})}{g(y^{n}|x^{n})}\right)\right]\cdot\exp(-\log(1/\delta))$
@@
$P\left(\sum_{n=1}^{N}\log\frac{b(y^{n}|x^{n})}{g(y^{n}|x^{n})}>c_{MLE}\log(\mathcal{N}_{\mathcal{F}_{g}}(1/N)/\delta)\right)\leq\delta/2,$
@@
$P\left(\sum_{n=1}^{N}\log\frac{\tilde{g}(y^{n}|x^{n})}{g(y^{n}|x^{n})}>c_{MLE}\log(\mathcal{N}_{\mathcal{F}_{g}}(1/N)/\delta)\right)\leq\delta/2.$
@@
$\mathbb{E}_{x\sim\mathcal{D}}\left[\|g(\cdot|x)-\tilde{g}(\cdot|x)\|_{1}^{2}\right]\leq\frac{c\log(\mathcal{N}_{\mathcal{F}_{g}}(1/N)/\delta)}{N},$ (4)
@@
The proof follows that of Liu et al. [2022], Proposition 14. ∎
This guarantees that the true reward function is within an interval around the MLE estimate with high probability.
@@
Let $\mathcal{E}_{R}$ denote the event $R\in\mathcal{R}$ and $\mathcal{E}_{T}$ denote the event $T\in\mathcal{T}$, $\mathcal{R}$ and $\mathcal{T}$ denote the respective confidence sets around the MLE. By Lemma A.1, events $\mathcal{E}_{R}$ and $\mathcal{E}_{T}$ have probability $21-\delta/$ 2 if we choose $\beta_{R}={c^{\prime}_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}/{N_{p}}$ and $\beta_{T}=c^{\prime}_{T}\log(H\mathcal{N}_{\mathcal{F}_{T}}(1/N_{o})/\delta)/N_{o}$, where $c^{\prime}_{R},c^{\prime}_{T}$ are universal constants.
@@ A.2 Model-based Pessimism and Uncertainty Penalties
$V_{T,{R}}^{\pi}-V_{\hat{T},{R}}^{\pi}$ $\leq\mathbb{E}_{\tau\sim d^{\pi}_{\hat{T}}}\left[\sum_{s_{j},a_{j}\in\tau}\|T(\cdot|s_{j},a_{j})-\hat{T}(\cdot|s_{j},a_{j})\|_{1}\right]\cdot R_{max}$
@@
$\leq\mathbb{E}_{s_{j},a_{j}\sim\pi,\hat{T}}\left[\|T(\cdot|s_{j},a_{j})-\hat{T}(\cdot|s_{j},a_{j})\|_{1}\cdot R_{max}\right]$
@@
$\leq\sum_{j=0}^{H-1}\mathbb{E}_{s_{j},a_{j}\sim\pi,\hat{T}}\left[\|T(\cdot|s_{j},a_{j})-\hat{T}(\cdot|s_{j},a_{j})\|_{1}\cdot R_{max}\right]$
$=\mathbb{E}_{\tau\sim d^{\pi}_{\hat{T}}}\left[\sum_{s_{j},a_{j}\in\tau}\|T(\cdot|s_{j},a_{j})-\hat{T}(\cdot|s_{j},a_{j})\cdot R_{max}\|_{1}\right]$
@@
where we have used the telescoping lemma (Lemma A.3), and where $u_{T}(\tau)=\sum_{(s,a)\in\tau}u_{T}(s,a)\geq\sum_{(s,a)\in\tau}\|\hat{T}(\cdot|s,a)-T(\cdot|s,a)\|_{1}\cdot R_{max}$ under event $\mathcal{E}_{T}$. ∎
@@
where we have used the fact that $|\hat{R}(\tau)-R(\tau)|\leq\sum_{s,a\in\tau}|\hat{R}(s,a)-R(s,a)|=\sum_{(s,a)\in\tau}u_{R}(s,a)=u_{R}(\tau)$ under event $\mathcal{E}_{R}$. ∎
@@
This justifies the overall objective considered in our pessimistic policy optimization procedure in Section 4.
@@ A.3 Suboptimality lower bound: a counterexample
Let $\pi^{*}_{offline}=\text{argmax}_{\pi\in\Pi}\min_{\tilde{T}\in\mathcal{T}}V_{\tilde{T},R}^{\pi}$ denote the optimal offline policy, which has access to the ground-truth reward function. In this section, we ask whether its suboptimality $\epsilon_{T}=V^{\pi^{*}}_{T,R}-V^{\pi^{*}_{offline}}_{T,R}$ is a lower bound for the suboptimality of our learned policy $\hat{\pi}^{*}$ after preference elicitation.
@@
Consider the MDP illustrated in Figure 4. Assume the following MLE estimate and uncertainty function for both the transition and reward models:
@@
$V^{a_{0}}_{\hat{T}_{inf},R}$ $=0.1\cdot 1+0.9\cdot 0=0.1$
$V^{a_{1}}_{\hat{T}_{inf},R}$ $=0.6\cdot 0+0.4\cdot 1=0.4$
Thus, we have: $\pi_{offline}^{*}(s_{0})=\text{argmax}_{a}V^{a}_{\hat{T}_{inf},R}=a_{1}$. The offline policy picks the suboptimal action since the worst-case returns of this action are lower than those estimated for $a_{0}$. Evaluating this policy in the real environment, we get $\epsilon_{T}=V^{\pi^{*}}_{T,R}-V^{\pi^{*}_{offline}}_{T,R}=0.6\cdot 0+0.4\cdot 1=0.4$.
We now estimate the optimal policy in the learned transition and reward model. Applying pessimism with respect to both models, we get an equal estimated value of 00 for both actions $a_{0}$ and $a_{1}$. If policy optimization converges to $\hat{\pi}^{*}=a_{0}$, we reach the suboptimality $V^{\pi^{*}}_{T,R}-V^{\hat{\pi}^{*}}_{T,R}=0.8\cdot 1+0.2\cdot 0=0.8>\epsilon_{T}$.
This example demonstrates that $\epsilon_{T}$ is not a lower bound for the suboptimality of $\hat{\pi}^{*}$, as policy $\hat{\pi}^{*}$ can achieve lower suboptimality than $\pi^{*}_{offline}$ if errors in transition and reward model estimation compensate each other.
@@ A.4.1 Suboptimality Decomposition
Recall that $\hat{T}_{inf},\hat{R}_{inf}=\text{argmin}_{\tilde{T}\in\mathcal{T},\tilde{R}\in\mathcal{R}}V_{\hat{T},\hat{R}}^{\pi}$ denote the pessimistic transition and reward models, such that $\hat{\pi}^{*}=\text{argmax}_{\pi\in\Pi}V_{\hat{T}_{inf},\hat{R}_{inf}}^{\pi}$. We have:
@@
$=(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{inf},\hat{R}_{inf}})-(V^{\hat{\pi}^{*}}_{T,R}-V^{{\pi}^{*}}_{\hat{T}_{inf},\hat{R}_{inf}})$
$\leq(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{inf},\hat{R}_{inf}})-(V^{\hat{\pi}^{*}}_{T,R}-V^{\hat{\pi}^{*}}_{\hat{T}_{inf},\hat{R}_{inf}})$
$\leq V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{inf},\hat{R}_{inf}},$ (5)
where we have first used the optimality of $\hat{\pi}^{*}$ (stating that $V^{\hat{\pi}^{*}}_{\hat{T}_{inf},\hat{R}_{inf}}\geq V^{\pi}_{\hat{T}_{inf},\hat{R}_{inf}}$, for all $\pi$) and then the pessimism principle (stating that $V^{\hat{\pi}^{*}}_{\hat{T}_{inf},\hat{R}_{inf}}\leq V^{\hat{\pi}^{*}}_{T,R}$).
@@
$V^{\pi^{*}}-V^{\hat{\pi}^{*}}$ $\leq\underbrace{(V^{{\pi}^{*}}_{T,\hat{R}_{inf}}-V^{\pi^{*}}_{\hat{T}_{inf},\hat{R}_{inf}})}_{\text{transition term}}+\underbrace{(V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{inf}})}_{\text{reward term}}$ (6)

### v2

@@ 7 Experimental results
We also measure how the coverage of the optimality of the dataset affects performance in our setting. In Figure 3(b), we vary the behavioral policy $\pi_{\beta}$ underlying the offline data, ranging from optimal (density ratio coefficient $=1$) to highly suboptimal (large density ratio coefficient). The concentrability terms $C_{T}$ and $C_{R}$ are challenging to measure as they require considering entire function classes, but we report the accuracy of the maximum likelihood estimate for both models in Appendix E. We observe that preference elicitation methods perform best when the data is close to optimal (with the exception of a fully optimal, non-diverse dataset making reward learning from preferences challenging). More preference samples are required if the observational dataset has poor coverage of the optimal policy (large $C_{T}(\mathcal{F}_{T},\pi^{*})$), as the transition and reward models become less accurate for the trajectory distribution of interest. We also validate this conclusion on HalfCheetah datasets of varying optimality in Appendix E.
@@
(b) As a function of dataset optimality.
Figure 3: Preference sample complexity $N_{p}$ as function of the properties of the observational data, to reach a suboptimality gap of $\epsilon=20$ over normalized environment returns (Star MDP). Mean and 95% confidence intervals over 6 experiments. $\times$ marks when the target suboptimality could not be achieved.
@@ 8 Conclusion
We provide theoretical guarantees on the sample complexity of both approaches, demonstrating that performance depends on how well the offline data covers the optimal policy. Empirical evaluations on various environments confirm the practical effectiveness of our algorithm, as Sim-OPRL consistently outperforms OPRL baselines in all settings.
@@ A.1 Maximum Likelihood and Confidence Intervals
$\beta=\frac{c_{\textrm{MLE}}}{N}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{g}}\left(\frac{1}{N}\right)\right),$
@@
Let $\bar{\mathcal{B}}$ be a $1/N$-bracket of $\mathcal{F}_{g}$ with |¯$|\bar{\mathcal{B}}|_{1}=\mathcal{N}_{\mathcal{F}_{g}}(1/N)$. Denote the set of all right brackets in $\bar{\mathcal{B}}$ by ~=$\tilde{\mathcal{B}}=\{b:\exists b^{\prime}s.t.[b^{\prime},b]\in\bar{\mathcal{B}}\}$. For $b\in\tilde{\mathcal{B}}$, we have:
@@
$P\left(\sum_{n=1}^{N}\log\frac{b(y^{n}|x^{n})}{g(y^{n}|x^{n})}>\log(\delta)\right)$ $\leq\mathbb{E}\left[\exp\left(\sum_{n=1}^{N}\log\frac{b(y^{n}|x^{n})}{g(y^{n}|x^{n})}\right)\right]\cdot\exp(-\log(1/\delta))$
@@
$P\left(\sum_{n=1}^{N}\log\frac{b(y^{n}|x^{n})}{g(y^{n}|x^{n})}>c_{MLE}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{g}}\left(\frac{1}{N}\right)\right)\right)\leq\delta/2,$
@@
$P\left(\sum_{n=1}^{N}\log\frac{\tilde{g}(y^{n}|x^{n})}{g(y^{n}|x^{n})}>c_{MLE}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{g}}\left(\frac{1}{N}\right)\right)\right)\leq\delta/2.$
@@
$\mathbb{E}_{x\sim\mathcal{D}}\left[\|g(\cdot|x)-\tilde{g}(\cdot|x)\|_{1}^{2}\right]\leq\frac{c}{N}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{g}}\left(\frac{1}{N}\right)\right),$ (6)
@@
The proof follows that of Liu et al. [2022a], Proposition 14. ∎
This guarantees that the true function is within an interval around the MLE estimate with high probability.
@@
Let $\mathcal{E}_{R}$ denote the event $R\in\mathcal{R}$ and $\mathcal{E}_{T}$ denote the event $T\in\mathcal{T}$, $\mathcal{R}$ and $\mathcal{T}$ denote the respective confidence sets around the MLE. By Lemma A.1, events $\mathcal{E}_{R}$ and $\mathcal{E}_{T}$ have probability $21-\delta/$ 2 if we choose $\beta_{R}={c^{\textrm{MLE}}_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}/{N_{p}}$ and $\beta_{T}=c^{\textrm{MLE}}_{T}\log(H\mathcal{N}_{\mathcal{F}_{T}}(1/N_{o})/\delta)/N_{o}$, where $c^{\textrm{MLE}}_{R},c^{\textrm{MLE}}_{T}$ are universal constants.
@@ A.2 Model-based Pessimism and Uncertainty Penalties
$V_{T,{R}}^{\pi}-V_{\hat{T},{R}}^{\pi}$ $\leq\mathbb{E}_{\tau\sim d^{\pi}_{\hat{T}}}\left[\sum_{s_{j},a_{j}\in\tau}\|T(\cdot|s_{j},a_{j})-\hat{T}(\cdot|s_{j},a_{j})\|_{1}\right]\cdot R_{\textrm{max}}$
@@
$\leq\mathbb{E}_{s_{j},a_{j}\sim\pi,\hat{T}}\left[\|T(\cdot|s_{j},a_{j})-\hat{T}(\cdot|s_{j},a_{j})\|_{1}\cdot R_{\textrm{max}}\right]$
@@
$\leq\sum_{j=0}^{H-1}\mathbb{E}_{s_{j},a_{j}\sim\pi,\hat{T}}\left[\|T(\cdot|s_{j},a_{j})-\hat{T}(\cdot|s_{j},a_{j})\|_{1}\cdot R_{\textrm{max}}\right]$
$=\mathbb{E}_{\tau\sim d^{\pi}_{\hat{T}}}\left[\sum_{s_{j},a_{j}\in\tau}\|T(\cdot|s_{j},a_{j})-\hat{T}(\cdot|s_{j},a_{j})\cdot R_{\textrm{max}}\|_{1}\right]$
@@
where we have used the telescoping lemma (Lemma A.3), and where $u_{T}(\tau)=\sum_{(s,a)\in\tau}u_{T}(s,a)\geq\sum_{(s,a)\in\tau}\|\hat{T}(\cdot|s,a)-T(\cdot|s,a)\|_{1}\cdot R_{\textrm{max}}$ under event $\mathcal{E}_{T}$. ∎
@@
where we have used the fact that $|\hat{R}(\tau)-R(\tau)|_{1}\leq\sum_{s,a\in\tau}|\hat{R}(s,a)-R(s,a)|_{1}=\sum_{(s,a)\in\tau}u_{R}(s,a)=u_{R}(\tau)$ under event $\mathcal{E}_{R}$. ∎
@@
This justifies the overall objective considered in our pessimistic policy optimization procedure in Section 6.3.
@@ A.3 Suboptimality lower bound: a counterexample
Let $\pi^{*}_{\textrm{offline}}=\text{argmax}_{\pi\in\Pi}\min_{\tilde{T}\in\mathcal{T}}V_{\tilde{T},R}^{\pi}$ denote the optimal offline policy, which has access to the ground-truth reward function. In this section, we ask whether its suboptimality $\epsilon_{T}=V^{\pi^{*}}_{T,R}-V^{\pi^{*}_{\textrm{offline}}}_{T,R}$ is a lower bound for the suboptimality of our learned policy $\hat{\pi}^{*}$ after preference elicitation.
@@
Consider the MDP illustrated in Figure 4. The value function for the optimal policy $\pi^{*}$, which consists of taking action $a_{0}$, is: $V^{\pi^{*}}_{T,R}=0.8\cdot 1+0.2\cdot 0=0.8$.
Now assume the following MLE estimate and uncertainty function for both the transition and reward models:
@@
$V^{a_{0}}_{\hat{T}_{\textrm{inf}},R}$ $=0.1\cdot 1+0.9\cdot 0=0.1$
$V^{a_{1}}_{\hat{T}_{\textrm{inf}},R}$ $=0.6\cdot 0+0.4\cdot 1=0.4$
Thus, we have: $\Pi_{\textrm{offline}}^{*}(s_{0})=\text{argmax}_{a}V^{a}_{\hat{T}_{\textrm{inf}},R}=a_{1}$. The offline policy picks the suboptimal action since the worst-case returns of this action are lower than those estimated for $a_{0}$. Evaluating this policy in the real environment, we get $V^{\pi^{*}_{\textrm{offline}}}_{T,R}=0.6\cdot 0+0.4\cdot 1=0.4$.
We now estimate the optimal policy in the learned transition and reward model. Applying pessimism with respect to both models, we get an equal estimated value of 00 for both actions $a_{0}$ and $a_{1}$. If policy optimization converges to $\hat{\pi}^{*}=a_{0}$, we achieve optimal performance with $V^{\hat{\pi}^{*}}_{T,R}=0.8>V^{\pi^{*}_{\textrm{offline}}}_{T,R}$.
This example demonstrates that $\epsilon_{T}=V^{\pi^{*}}_{T,R}-V^{\pi^{*}_{\textrm{offline}}}_{T,R}=0.4$ is not a lower bound for the suboptimality of $\hat{\pi}^{*}$, as policy $\hat{\pi}^{*}$ can achieve better performance (even optimal performance in this case) if errors in transition and reward model estimation compensate each other.
@@ A.4.1 Suboptimality Decomposition
Recall that $\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}=\text{argmin}_{\tilde{T}\in\mathcal{T},\tilde{R}\in\mathcal{R}}V_{\hat{T},\hat{R}}^{\pi}$ denote the pessimistic transition and reward models, such that $\hat{\pi}^{*}=\text{argmax}_{\pi\in\Pi}V_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}}^{\pi}$. We have:
@@
$=(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})-(V^{\hat{\pi}^{*}}_{T,R}-V^{{\pi}^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})$
$\leq(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})-(V^{\hat{\pi}^{*}}_{T,R}-V^{\hat{\pi}^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})$
$\leq V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}},$ (7)
where we have first used the optimality of $\hat{\pi}^{*}$ (stating that $V^{\hat{\pi}^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}}\geq V^{\pi}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}}$, for all $\pi$) and then the pessimism principle (stating that $V^{\hat{\pi}^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}}\leq V^{\hat{\pi}^{*}}_{T,R}$).
@@
$V^{\pi^{*}}-V^{\hat{\pi}^{*}}$ $\leq\underbrace{(V^{{\pi}^{*}}_{T,\hat{R}_{\textrm{inf}}}-V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})}_{\text{transition term}}+\underbrace{(V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{\textrm{inf}}})}_{\text{reward term}}$ (8)